# Visual Perturbation and Adaptive Hard Negative Contrastive Learning for Compositional Reasoning in Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) are essential for multimodal tasks, especially compositional reasoning (CR) tasks, which require distinguishing fine-grained semantic differences between visual and textual embeddings. However, existing methods primarily fine-tune the model by generating text-based hard negative samples, neglecting the importance of image-based negative samples, which results in insufficient training of the visual encoder and ultimately impacts the overall performance of the model. Moreover, negative samples are typically treated uniformly, without considering their difficulty levels, and the alignment of positive samples is insufficient, which leads to challenges in aligning difficult sample pairs. To address these issues, we propose Adaptive Hard Negative Perturbation Learning (AHNPL). AHNPL translates text-based hard negatives into the visual domain to generate semantically disturbed image-based negatives for training the model, thereby enhancing its overall performance. AHNPL also introduces a contrastive learning approach using a multimodal hard negative loss to improve the model’s discrimination of hard negatives within each modality and a dynamic margin loss that adjusts the contrastive margin according to sample difficulty to enhance the distinction of challenging sample pairs. Experiments on three public datasets demonstrate that our method effectively boosts VLMs’ performance on complex CR tasks. The source code is available at https://github.com/nynu-BDAI/AHNPL.
@@ 1 Introduction
In recent years, Visual Language Models (VLMs) have achieved significant zero-shot performance advancements in multimodal tasks such as retrieval Huang et al. (2017, 2025), classification Metzen et al. (2023); Novack et al. (2023) and segmentation Xu et al. (2022); Lüddecke and Ecker (2022), establishing themselves as a foundation for multimodal tasks. However, despite their impressive performance, VLMs still face notable challenges in handling compositional reasoning (CR) tasks Doveh et al. (2023a); Thrush et al. (2022). CR tasks demand that models accurately understand and reason about fine-grained relationships among objects, attributes, and actions in images or texts. For instance, they must be able to distinguish sentences differing only in word order or accurately identify objects in images that have similar colors but different attributes. Such tasks go beyond the need for coarse-grained semantic understanding, requiring precise reasoning in complex scenarios.
@@
However, current methods for incorporating hard negative samples into contrastive learning face several issues. First, some methods focus solely on hard negatives in the text modality, overlooking hard negatives in the visual modality Huang et al. (2024), resulting in insufficient training of the image encoder and thereby affecting the overall model performance. Second, these methods typically treat all negative samples equally, without considering their difficulty levels Zhang et al. (2024c, a). This prevents the model from effectively capturing subtle differences in hard negatives during training, thus hindering its ability to learn fine-grained distinctions. Additionally, these methods often neglect the adjustment of alignment for positive samples, which may cause the model to align only simple positive pairs while struggling to align more challenging positive pairs.
To address these issues, we propose Adaptive Hard Negative Perturbation Learning (AHNPL) to enhance VLMs’ performance in CR tasks. Our AHNPL method explicitly maps the subtle semantic variations of text-based hard negatives into the visual space, generating corresponding image embeddings for these negative texts, thereby improving the model’s ability to discern subtle differences in the visual modality, which ultimately enhances the overall performance of the model. Moreover, AHNPL introduces an dynamic hard negative contrastive learning, comprising a multimodal hard negative loss and a dynamic margin contrastive loss. The multimodal hard negative loss reduces the similarity between negative and positive samples across both text and visual modalities, improving the model’s ability to distinguish hard negatives. The dynamic margin contrastive loss adjusts the contrastive margin based on sample complexity, enabling the model to focus more on challenging negatives and positives, which improves the alignment quality of challenging sample pairs.
@@ 2 Related Work
Contrastive Learning    The objective of contrastive representation learning is to learn representations that are close to each other for similar samples and distant from each other for dissimilar samples. As one of the representative examples, CLIP Radford et al. (2021) has become a milestone in this field. CLIP is a Transformer-based Vaswani et al. (2017) model that consists of an image encoder and a text encoder, which are trained simultaneously. The objective is to maximize the cosine similarity of the image and text embeddings from the correct image-text pairs and to minimize the similarity between the incorrect pairs. A batch of N training samples (i.e., matching image-text pairs) results in a similarity matrix for each image-text combination. The main diagonal indicates the correct pair matches; the remaining entries correspond to negative entries. The InfoNCE loss is applied to the N × N similarity score matrix, and through contrastive learning, the image and text encoders are more effectively aligned across the two modalities, thus enhancing the performance of VLMs in multimodal tasks.
@@
Figure 2: The propose Adaptive Hard Negative Perturbation Learning (AHNPL). (a) Performing visual perturbation by calculating the semantic shift in text to generate a difference vector. (b) The dynamic hard negative contrastive learning method makes targeted adjustments to the strategies for alignment of positive samples and the distinction of hard negative samples.
@@
Compositional Reasoning   CR in vision and language tasks evaluates a model’s ability to understand and manipulate complex ideas by breaking them into simpler components and recombining them in new ways. While VLMs excel at handling multimodal data, research shows they struggle with analyzing complex information, such as object relationships. Thrush et al. Thrush et al. (2022) first identified this issue, demonstrating that VLMs often fail to distinguish semantic differences caused by changes in word order, with performance sometimes no better than random guessing. This type of task, which tests a model’s ability to understand fine-grained semantic structures, is referred to as CR. Many researchers have proposed methods of generating negative samples to fine-tune models to improve their performance in CR tasks. For example, Doveh et al. Doveh et al. (2023a) introduced DAC, which fine-tunes the model to improve its performance in CR tasks by using a Large Language Model (LLM) to generate texts of different scenes. Zhang et al. Zhang et al. (2024b) proposed CE-CLIP, which refines contrastive objectives with diverse negative samples to enhance semantic understanding. While effective, we believe current methods haven’t fully explored hard negative mining, as they overlook the importance of image-based negative samples and fail to adjust the alignment process of positive sample pairs. To address this issue, we propose AHNPL, which translates text-based hard negatives into the visual domain, generating semantically disturbed image-based negatives, and dynamically adjusts the training process based on the similarity of each positive sample pair.
@@ 3 Methodology
In this section, we provide a detailed introduction to our proposed AHNPL. Fig. 2 shows an overview of the entire pipeline.
@@ 3.2 Visual Perturbation with Semantic Shift
In the previous section, we identified keywords in the captions through part-of-speech parsing and constructed textual negative samples. Furthermore, we design a method for generating visual negative samples to train the model’s image encoder, thereby enhancing the model’s overall performance.
Existing negative sample generation methods often overly focus on textual negative samples, neglecting the role of visual negative samples, which leads to insufficient training of the image encoder and subsequently affects the overall model performance. To solve this problem, we propose a visual perturbation with semantic shift method for generating image negative samples. Specifically, we first capture semantic changes by calculating the deviation embedding between the original text and the negative text. Given the original text $T_{{orig}}$ and the negative text $T_{{neg}}$, we use a pretrained CLIP model to extract their high-dimensional embedding vectors $e_{T_{{orig}}}$ and $e_{T_{{neg}}}$, respectively. These vectors represent the semantic information of the texts within the multimodal embedding space of the CLIP model. The deviation embedding $\Delta_{e}$ is calculated as follows:
@@ 3.3 Dynamic Hard Negative Contrastive Learning
Contrastive Loss   Our method uses a contrastive loss, which is applied to a text and image pair $(T,I)$ as input and consists of two components: (i) an image encoder $e_{I}=f_{v}(I)$; (ii) a text encoder $e_{T}=f_{t}(T)$. In this setting, the text-to-image similarity score is computed as:
@@
Multimodal Hard Negative Loss   To enhance the model’s ability to distinguish negative samples, we specially introduce a negative sample loss term. By simultaneously handling negative samples from both visual and textual perspectives, we help the model more accurately distinguish these challenging negative samples, thereby improving the effectiveness of contrastive learning.
@@
Table 1: Results (%) on ARO and VALSE. Our proposed AHNPL achieved new state-of-the-art (SOTA) results on the VALSE benchmark.
Dynamic Margin Contrastive Loss  In contrastive learning, the alignment of positive samples and the distinction of negative samples are key to improving model performance. Existing methods typically use a fixed margin (Margin Loss) threshold to adjust the similarity between positive and negative samples, but they still perform poorly when handling samples of varying difficulty.
When it comes to positive samples, such fixed margin approaches actually hold the assumption that all samples have the same level of difficulty. As a result, when handling challenging samples, the model tends to treat them the same as simple samples, making it difficult to capture the subtle semantic relationships in the complexity, ultimately reducing the model’s ability to align positive samples effectively. For negative samples, these methods generally introduce them into the contrastive learning framework without considering the variations in their difficulty. Although hard negatives are semantically similar to positive samples, they are characterized by subtle semantic differences. This subtle semantic difference makes it challenging for the model to effectively distinguish between positive samples and hard negatives, thereby limiting the model’s performance.
@@
where $M_{n}^{t}$ is the adaptive threshold calculated for each textual hard negative $T_{n}$ in the training step $t$. This threshold is calculated as follows:
@@ 4 Experiments
In this section, we describe the experimental setup, including the datasets used, baseline models for comparison, and the evaluation environment. We compare our method with several existing approaches, and the results show effective performance improvements across multiple benchmark datasets.
@@ 4.2 Dataset
Training Datasets. The CLIP model is originally pretrained on 400 million image-text pairs sourced from the web, and we continue to fine-tune the model based on this pretraining. In our experiments, we use the widely adopted cross-modal text-image retrieval dataset MSCOCO Lin et al. (2014). MSCOCO is chosen for its rich annotations of diverse objects, attributes, and relationships, which enhance the model’s semantic understanding in complex multimodal tasks.
Evaluation Dataset. We evaluated our method on several vision-language compositional benchmarks: ARO, VALSE Parcalabescu et al. (2022), and SugarCrepe Hsieh et al. (2023) (bias-mitigated version of CREPE Yu et al. (2023)). Each test example in these datasets includes an image along with a corresponding correct description and a modified incorrect description. The model’s task is to determine which description.
  • ARO Yuksekgonul et al. (2022). This is a CR benchmark that includes positive or negative captions and evaluates sensitivity to word order. Word-order negative sentences are created by reordering words, which changes sentence semantics in attributes, relationships, and word order meaning. ARO is constructed from images and captions from COCO, Flick30K Young et al. (2014), and Visual Genome Krishna et al. (2017).
@@
Figure 4: Predictions of CLIP and AHNPL on SugarCrepe tasks: Swap Object, Swap Attribute, Replace Relationship and Replace Attribute. The score represents the similarity score between the caption and the corresponding image as assessed by CLIP/AHNPL. The model selects the caption with the higher similarity score as the correct one.
@@ 4.3 Overall Results
We present the evaluation results for the ARO and VALSE benchmarks in Tab. 1, and Tab. 2 shows the evaluation results on the SugarCrepe benchmark. As a widely used evaluation benchmark in this field, VALSE provides a solid basis for comparing our model against various mainstream models. Under the same training conditions using the CE-CLIP dataset, our AHNPL demonstrated effective improvements over all methods that utilized hard negatives, achieving excellent results across all benchmarks.
@@
80.1 71.9 72.7 87.5 78.1
Ours 83.8 77.0 75.9 93.6 82.3
Table 3: Ablation of losses, where negatives represents image-text contrastive with additional hard negatives.
As shown in Tab. 1, our method achieved new state-of-the-art (SOTA) results on the VALSE dataset, with an average matching score of 75.9%, outperforming the current SOTA model CE-CLIP by 3.4%. Other models, such as NegCLIP, also performed well with a score of 71.6%, but our model still led by 4.3%, particularly excelling in the quantifiers and clean subtasks, where it outperformed CE-CLIP by 5.4% and 3.8%, respectively. On the ARO benchmark, AHNPL outperformed CE-CLIP in the ARO-Relation task, indicating an improvement in relationship understanding. In the ARO-Attribute task, AHNPL slightly surpassed CE-CLIP, suggesting that our model is comparable to the current SOTA model in attribute understanding.
@@ 4.5 Case Study
We present several case studies illustrating the performance of CLIP and AHNPL across four subtasks of the SugarCrepe benchmark, as shown in Fig. 4. SugarCrepe uses LLMs to generate fluent captions with common-sense, challenging VLMs to distinguish negative captions effectively. In the “Swap Object” task, where models must understand object relationships (e.g., “A painting of a vase with a sunflower on a table” vs. “A painting of a sunflower with a vase on a table”), AHNPL outperforms CLIP. In the “Swap Attribute” task requiring accurate identification of object attributes (e.g., “Four surfers are trying to catch a wave as they stand” vs. “Trying to catch four waves, surfers stand”), CLIP struggles, whereas AHNPL consistently selects the correct caption. For the “Replace Relationship” and “Replace Attribute” tasks, which involve subtle distinctions (e.g., “A man is flying a kite at the beach” vs. “A man is holding a kite at the beach” or “Three teddy bears laying in bed under the covers” vs. “A single teddy bear laying in bed under the covers”), AHNPL effectively handles these nuances through negative caption contrastive learning, outperforming CLIP.

### v2

@@ Abstract
Vision-Language Models (VLMs) are essential for multimodal tasks, especially compositional reasoning (CR) tasks, which require distinguishing fine-grained semantic differences between visual and textual embeddings. However, existing methods primarily fine-tune the model by generating text-based hard negative samples, neglecting the importance of image-based negative samples, which results in insufficient training of the image encoder and ultimately impacts the overall performance of the model. Moreover, negative samples are typically treated uniformly, without considering their difficulty levels, and the alignment of positive samples is insufficient, which leads to challenges in aligning difficult sample pairs. To address these issues, we propose Adaptive Hard Negative Perturbation Learning (AHNPL). AHNPL translates text-based hard negatives into the visual domain to generate semantically disturbed image-based negatives for training the model, thereby enhancing its overall performance. AHNPL also introduces a contrastive learning approach using a multimodal hard negative loss to improve the model’s discrimination of hard negatives within each modality and a dynamic margin loss that adjusts the contrastive margin according to sample difficulty to enhance the distinction of challenging sample pairs. Experiments on three public datasets demonstrate that our method effectively boosts VLMs’ performance on complex CR tasks. The source code is available at https://github.com/nynu-BDAI/AHNPL.
@@ 1 Introduction
In recent years, Vision-Language Models (VLMs) have achieved significant zero-shot performance advancements in multimodal tasks such as retrieval Huang et al. (2017, 2025), classification Metzen et al. (2023); Novack et al. (2023) and segmentation Xu et al. (2022); Lüddecke and Ecker (2022), establishing themselves as a foundation for multimodal tasks. However, despite their impressive performance, VLMs still face notable challenges in handling compositional reasoning (CR) tasks Doveh et al. (2023a); Thrush et al. (2022). CR tasks demand that models accurately understand and reason about fine-grained relationships among objects, attributes, and actions in images or texts. For instance, they must be able to distinguish sentences differing only in word order or accurately identify objects in images that have similar colors but different attributes. Such tasks go beyond the need for coarse-grained semantic understanding, requiring precise reasoning in complex scenarios.
@@
However, current methods for incorporating hard negative samples into contrastive learning face several issues. First, some methods focus solely on hard negatives in the text modality, overlooking equally important hard negatives in the visual modality Huang et al. (2024), resulting in insufficient training of the image encoder and thereby affecting the overall model performance. Second, these methods typically treat all negative samples equally, without considering their varying difficulty levels Zhang et al. (2024c, a). This prevents the model from effectively capturing subtle differences in hard negatives during training, thus hindering its ability to learn fine-grained distinctions. Additionally, these methods often neglect the adjustment of alignment for positive samples, which may cause the model to align only simple positive pairs while struggling to align more challenging ones.
To address these issues, we propose Adaptive Hard Negative Perturbation Learning (AHNPL) to enhance VLMs’ performance in CR tasks. Our AHNPL method explicitly maps the subtle semantic variations of text-based hard negatives into the visual space, generating corresponding image embeddings for these negative texts, thereby improving the model’s ability to discern subtle differences in the visual modality, which ultimately enhances the overall performance of the model. Moreover, AHNPL introduces dynamic hard negative contrastive learning, comprising a multimodal hard negative loss and a dynamic margin contrastive loss. The multimodal hard negative loss reduces the similarity between negative and positive samples across both text and visual modalities, improving the model’s ability to distinguish hard negatives. The dynamic margin contrastive loss adjusts the contrastive margin based on sample complexity, enabling the model to focus more on challenging negatives and positives, which improves the alignment quality of challenging sample pairs.
@@ 2 Related Work
The objective of contrastive representation learning is to learn representations that are close to each other for similar samples and distant from each other for dissimilar samples. As one of the representative examples, CLIP Radford et al. (2021) has become a milestone in this field. CLIP is a Transformer-based Vaswani et al. (2017) model that consists of an image encoder and a text encoder, which are trained simultaneously. The objective is to maximize the cosine similarity of the image and text embeddings from the correct image-text pairs and to minimize the similarity between the incorrect pairs. A batch of N training samples (i.e., matching image-text pairs) results in a similarity matrix for each image-text combination. The main diagonal indicates the correct pair matches; the remaining entries correspond to negative entries. The InfoNCE loss is applied to the N × N similarity score matrix, and through contrastive learning, the image and text encoders are more effectively aligned across the two modalities, thus enhancing the performance of VLMs in multimodal tasks.
@@
Figure 2: The proposed Adaptive Hard Negative Perturbation Learning (AHNPL). (a) Performing visual perturbation by computing the semantic shift in text to generate a deviation embedding. (b) The dynamic hard negative contrastive learning method makes targeted adjustments to the strategies for alignment of positive samples and the distinction of hard negative samples.
Compositional Reasoning.
@@
CR in vision and language tasks evaluates a model’s ability to understand and manipulate complex ideas by breaking them into simpler components and recombining them in new ways. While VLMs excel at handling multimodal data, research shows they struggle with analyzing complex information, such as object relationships. Thrush et al. Thrush et al. (2022) first identified this issue, demonstrating that VLMs often fail to distinguish semantic differences caused by changes in word order, with performance sometimes no better than random guessing. This type of task, which tests a model’s ability to understand fine-grained semantic structures, is referred to as CR. Many researchers have proposed methods of generating negative samples to fine-tune models to improve their performance in CR tasks. For example, Doveh et al. Doveh et al. (2023a) introduced DAC, which fine-tunes the model to improve its performance in CR tasks by using a Large Language Model (LLM) to generate texts of different scenes. Zhang et al. Zhang et al. (2024b) proposed CE-CLIP, which refines contrastive objectives with diverse negative samples to enhance semantic understanding. While effective, we believe current methods haven’t fully explored hard negative mining, as they overlook the importance of image-based negative samples and fail to adjust the alignment process of positive sample pairs. To address this issue, we propose AHNPL, which translates text-based hard negatives into the visual domain, generating semantically disturbed image-based negatives, and dynamically adjusts the training process based on the similarity of each positive sample pair.
@@ 3 Methodology
In this section, we provide a detailed introduction to our proposed AHNPL method. Fig. 2 shows an overview of the entire pipeline.
@@ 3.2 Visual Perturbation with Semantic Shift
In the previous section, we identify keywords in the captions through part-of-speech parsing and construct textual negative samples. Furthermore, we design a method for generating visual negative samples to train the model’s image encoder, thereby enhancing the model’s overall performance.
Existing negative sample generation methods often overly focus on textual negative samples, neglecting the role of visual negative samples, which leads to insufficient training of the image encoder and subsequently affects the overall model performance. To solve this problem, we propose a visual perturbation with semantic shift method for generating image negative samples. Specifically, we first capture semantic changes by computing the deviation embedding between the original text and the negative text. Given the original text $T_{{orig}}$ and the negative text $T_{{neg}}$, we use a pretrained CLIP model to extract their high-dimensional embedding vectors $e_{T_{{orig}}}$ and $e_{T_{{neg}}}$, respectively. These vectors represent the semantic information of the texts within the multimodal embedding space of the CLIP model. The deviation embedding $\Delta_{e}$ is computed as follows:
@@ 3.3 Dynamic Hard Negative Contrastive Learning
Contrastive Loss
Our method uses a contrastive loss, which is applied to a text and image pair $(T,I)$ as input and consists of two components: (i) an image encoder $e_{I}=f_{v}(I)$; (ii) a text encoder $e_{T}=f_{t}(T)$. In this setting, the text-to-image similarity score is computed as:
@@
Multimodal Hard Negative Loss.
To enhance the model’s ability to distinguish negative samples, we specially introduce a negative sample loss term. By simultaneously handling negative samples from both visual and textual perspectives, we help the model more accurately distinguish these challenging negative samples, thereby improving the effectiveness of contrastive learning.
@@
Table 1: Results (%) on ARO and VALSE. Our proposed AHNPL achieves new state-of-the-art (SOTA) results on the VALSE benchmark.
Dynamic Margin Contrastive Loss.
In contrastive learning, the alignment of positive samples and the distinction of negative samples are key to improving model performance. Existing methods typically use a fixed margin (Margin Loss) threshold to adjust the similarity between positive and negative samples, but they still perform poorly when handling samples of varying difficulty.
When it comes to positive samples, such fixed margin approaches actually hold the assumption that all samples have the same level of difficulty. As a result, when handling challenging samples, the model tends to treat them the same as simple samples, making it difficult to capture the subtle semantic relationships, ultimately reducing the model’s ability to align positive samples effectively. For negative samples, these methods generally introduce them into the contrastive learning framework without considering the variations in their difficulty. Although hard negatives are semantically similar to positive samples, they are characterized by subtle semantic differences. This subtle semantic difference makes it challenging for the model to effectively distinguish between positive samples and hard negatives, thereby limiting the model’s performance.
@@
where $M_{n}^{t}$ is the adaptive threshold computed for each textual hard negative $T_{n}$ in the training step $t$. This threshold is computed as follows:
@@ 4.2 Dataset
Training Datasets.
The CLIP model is originally pretrained on 400 million image-text pairs sourced from the web, and we continue to fine-tune the model based on this pretraining. In our experiments, we use the widely adopted cross-modal text-image retrieval dataset MSCOCO Lin et al. (2014). MSCOCO is chosen for its rich annotations of diverse objects, attributes, and relationships, which enhance the model’s semantic understanding in multimodal tasks.
Evaluation Dataset.
We evaluate our method on several vision-language compositional benchmarks: ARO, VALSE Parcalabescu et al. (2022), and SugarCrepe Hsieh et al. (2023) (a bias-mitigated version of CREPE Yu et al. (2023)). Each test example in these datasets includes an image along with a corresponding correct description and a modified incorrect description. The model’s task is to determine which description is correct.
  • ARO Yuksekgonul et al. (2022). This is a CR benchmark that includes positive or negative captions and evaluates sensitivity to word order. Word-order negative sentences are created by reordering words, which changes sentence semantics in attributes, relationships, and word order meaning.
@@ 4.3 Overall Results
We present the evaluation results for the ARO and VALSE benchmarks in Tab. 1, and Tab. 2 shows the evaluation results on the SugarCrepe benchmark. As a widely used evaluation benchmark in this field, VALSE provides a solid basis for comparing our model against various mainstream models. Our AHNPL demonstrates effective improvements over all methods that utilize hard negatives, achieving excellent results across all benchmarks.
@@
80.1 71.9 72.7 85.5 77.6
Ours 83.8 77.0 75.9 86.5 80.8
Table 3: Ablation of losses, where negatives represent image-text contrastive with additional hard negatives.
As shown in Tab. 1, our method achieves new state-of-the-art (SOTA) results on the VALSE dataset, with an average matching score of 75.9%, outperforming the current SOTA model CE-CLIP by 3.4%. Other models, such as NegCLIP, also perform well with a score of 71.6%, but our model still leads by 4.3%, particularly excelling in the quantifiers and clean subtasks, where it outperforms CE-CLIP by 5.4% and 3.8%, respectively. On the ARO benchmark, AHNPL outperforms CE-CLIP in the ARO-Relation task, indicating an improvement in relationship understanding. In the ARO-Attribute task, AHNPL slightly surpasses CE-CLIP, suggesting that our model is comparable to the current SOTA model in attribute understanding.
@@ 4.4 Ablation Study
Table 4: Predictions of CLIP and AHNPL on SugarCrepe tasks: Swap Object, Swap Attribute, Replace Relationship and Replace Attribute. The score represents the similarity score between the caption and the corresponding image as assessed by CLIP/AHNPL. The model selects the caption with the higher similarity score as the correct one.
@@ 4.5 Case Study
We present several case studies illustrating the performance of CLIP and AHNPL across four subtasks of the SugarCrepe benchmark, as shown in Tab. 4. SugarCrepe uses LLMs to generate fluent captions with common sense, thereby challenging VLMs to distinguish negative captions effectively. In the “Swap Object” task, where models must understand object relationships (e.g., “A painting of a vase with a sunflower on a table” vs. “A painting of a sunflower with a vase on a table”), AHNPL outperforms CLIP. In the “Swap Attribute” task requiring accurate identification of object attributes (e.g., “Four surfers are trying to catch a wave as they stand” vs. “Trying to catch four waves, surfers stand”), CLIP struggles, whereas AHNPL consistently selects the correct caption. For the “Replace Relationship” and “Replace Attribute” tasks, which involve subtle distinctions (e.g., “A man is flying a kite at the beach” vs. “A man is holding a kite at the beach” or “Three teddy bears laying in bed under the covers” vs. “A single teddy bear laying in bed under the covers”), AHNPL effectively handles these nuances through negative caption contrastive learning, outperforming CLIP.